Project

General

Profile

Actions

2015-07-28 Physics coordinator report

Beamline
  • VB has been asked to report at the analysis meeting on Thursday
Tracker alignment (no field)
  • Chris Hunt has produced residuals; compare well with MC
  • Note still no PID available from MAUS; analysis group working on dumb cuts solution for now
Tracker alignment (field)
  • Rogers has produced residuals; but some problem with the raw data - looks like cabling issue. Tracker group to generate new cabling and calibration files and data will need to be reprocessed. Highlights need for better online plots.
2015/02
  • First planning meeting tomorrow.
MICE Demonstration of Ionisation Cooling
  • The 140 MeV/c optics did not look so good, with poor transmission; JB and JPa are looping back through it.
  • Seek to ramp Strathclyde on doing alignment and tolerances and then looking at alternate optics solutions.

Updated by Rogers, Chris over 7 years ago ยท 5 revisions